OpenStudy (anonymous):

Carlos solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x + 10 + 4x = 10 5x + 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 He should have substituted 5 + 2x He did not distribute correctly He combined like terms incorrectly, it should have been 4x instead of 5x He made no mistake

OpenStudy (anonymous):

x − 2(5 − 2x) = 10 x + 10 + 4x = 10 look at those two lines what's wrong with it ?

OpenStudy (anonymous):

the minus changed to a + in the beginning of the equation, and im not sure

OpenStudy (anonymous):

good.. thats the problem

OpenStudy (anonymous):

so it should be -4x

OpenStudy (anonymous):

no.. -2(5-2x) = -10 +4x

OpenStudy (anonymous):

oh